How many orbitals are present in p- subshell ? Please answer fast

Grade:11

1 Answers

The Hounds Of Justice
5 years ago
The number of orbitals in a subshell are equal to = 2l+1
For p-subshell,’l’ value is = 1
so 2l+1= 2(1)+1 = 3
Sp, there are three orbitals in p-subshell